Grade 11Mechanics

A ro rocket is fired vertically and since with a constant acceleration of 20 metre per second square for 1 minute its fuel is then all use and it continued as a free fall vary what is the maximum altitude reached what is the total time lapse from the take off until rocket strikes Earth

Ieemaan Ashraf 310
5 Years agoGrade 11

1 Answer

Harshit Singh
5 Years ago
Dear Student

S=ut-1​/2 at2
S=1​/2(10)(30)2
=4500m
now
v2−u2=2as
v=300m/s​
now after fuel is consumed
a=−10
v=0
v=u−gt
t=30s
total time=30+30=60s​



Thanks